Calculate the savings you would have in the account at the end of 12 months if you made the following deposits. The banks pays 8% interest per year, compounded quarterly and it pays simple interest on any inter period deposits. End of Month 2 4 6 8 9 10 Deposits 250 75 65 105 95 95
Answer #1

Deposit made on 2nd month=250

Effective Monthly interest rate={[(1+(8%/4))^4-1]/12}*100=0.69%

At the end of 2nd month  Deposit would be =250+(0.0069)*10*250-267.25

At the end of 4th month  Deposit would be =75+(0.0069)*8*75=79.14

At the end of 6th month  Deposit would be =65+(0.0069)*6*65=67.69

At the end of 8th month  Deposit would be =105+(0.0069)*4*105=107.89

At the end of 9th month  Deposit would be =95+(0.0069)*95*3=96.97

At the end of 10th month  Deposit would be =95+(0.0069*95)*2=96.311

Hence sum would give the available deposit at the end of 12 months=267.25+79.14+67.69+107.89+96.97+96.31=$715.25

Total saving=$715.25